# Towards a Unified Copernicus Foundation Model for Earth Vision

### v1

@@ 5 Copernicus-Bench
Metric Supervised Supervised Random SoftCon [64] CROMA [24] DOFA [69] Copernicus-FM Backbone ViT-S/16 ViT-B/16 ViT-B/16 ViT-B/14 ViT-B/8 ViT-B/16 ViT-B/16* Modality S1/S2 S1+S2 All (spectral) All Cloud-S2 mIoU 64.2 $\pm$ 0.9 59.4 $\pm$ 1.0 60.4 $\pm$ 0.2 66.9 $\pm$ 0.3 65.0 $\pm$ 0.2 65.0 $\pm$ 0.2 66.7 $\pm$ 0.1 Cloud-S3 mIoU 61.7 $\pm$ 0.7 63.0 $\pm$ 0.8 60.9 $\pm$ 0.0 58.2 $\pm$ 0.1 62.0 $\pm$ 0.7 EuroSAT-S1 OA 81.7 $\pm$ 0.7 81.5 $\pm$ 0.9 75.4 $\pm$ 0.4 83.6 $\pm$ 0.1 83.9 $\pm$ 0.1 81.7 $\pm$ 0.1 87.2 $\pm$ 0.1 EuroSAT-S2 OA 97.5 $\pm$ 0.0 97.6 $\pm$ 0.1 92.5 $\pm$ 0.1 96.7 $\pm$ 0.0 97.0 $\pm$ 0.1 97.2 $\pm$ 0.1 97.9 $\pm$ 0.1 BigEarthNet-S1 mAP 78.1 $\pm$ 0.6 81.2 $\pm$ 0.5 66.1 $\pm$ 0.1 81.6 $\pm$ 0.0 72.8 $\pm$ 0.0 74.3 $\pm$ 0.0 83.3 $\pm$ 0.0 BigEarthNet-S2 mAP 83.6 $\pm$ 0.4 86.4 $\pm$ 0.4 73.3 $\pm$ 0.1 86.1 $\pm$ 0.0 78.8 $\pm$ 0.0 79.7 $\pm$ 0.0 84.6 $\pm$ 0.0 LC100Cls-S3 mAP 91.3 $\pm$ 0.3 91.4 $\pm$ 0.5 88.9 $\pm$ 0.1 89.5 $\pm$ 0.0 93.3 $\pm$ 0.4 DFC2020-S1 mIoU 49.9 $\pm$ 0.4 50.8 $\pm$ 0.5 45.4 $\pm$ 0.1 52.8 $\pm$ 0.6 52.7 $\pm$ 0.1 49.7 $\pm$ 0.1 52.4 $\pm$ 0.1 DFC2020-S2 mIoU 65.3 $\pm$ 0.6 66.2 $\pm$ 0.7 62.3 $\pm$ 0.0 64.1 $\pm$ 0.3 66.5 $\pm$ 0.0 61.8 $\pm$ 0.1 64.5 $\pm$ 0.1 LC100Seg-S3 mIoU 20.1 $\pm$ 0.4 19.3 $\pm$ 0.5 18.2 $\pm$ 0.1 16.5 $\pm$ 0.1 24.1 $\pm$ 0.0 Flood-S1 mIoU 78.0 $\pm$ 0.1 78.3 $\pm$ 0.3 75.1 $\pm$ 0.1 77.2 $\pm$ 0.1 77.4 $\pm$ 0.1 76.0 $\pm$ 0.1 77.7 $\pm$ 0.0 LCZ-S2 OA 86.6 $\pm$ 0.7 85.3 $\pm$ 0.8 77.4 $\pm$ 0.1 83.6 $\pm$ 0.2 84.1 $\pm$ 0.0 83.0 $\pm$ 0.3 84.4 $\pm$ 0.0 Biomass-S3 RMSE $\downarrow$ 68.1 $\pm$ 0.3 68.3 $\pm$ 0.4 68.7 $\pm$ 0.5 74.1 $\pm$ 0.1 66.3 $\pm$ 0.1 AQ-NO2-S5P RMSE $\downarrow$ 3.4 $\pm$ 0.0 3.4 $\pm$ 0.0 3.4 $\pm$ 0.0 3.3 $\pm$ 0.0 2.8 $\pm$ 0.0 AQ-O3-S5P RMSE $\downarrow$ 1781.3 $\pm$ 29.8 1766.8 $\pm$ 22.1 1741.6 $\pm$ 11.5 1755.6 $\pm$ 19.8 789.4 $\pm$ 2.6

### v2

@@ 5 Copernicus-Bench
Metric Supervised Supervised Random SoftCon [64] CROMA [24] DOFA [69] Copernicus-FM Backbone ViT-S/16 ViT-B/16 ViT-B/16 ViT-B/14 ViT-B/8 ViT-B/16 ViT-B/16* Modality S1/S2 S1+S2 All (spectral) All Cloud-S2 mIoU 64.2 $\pm$ 0.9 59.4 $\pm$ 1.0 60.4 $\pm$ 0.2 66.9 $\pm$ 0.3 65.0 $\pm$ 0.2 65.0 $\pm$ 0.2 66.7 $\pm$ 0.1 Cloud-S3 mIoU 61.7 $\pm$ 0.7 63.0 $\pm$ 0.8 60.9 $\pm$ 0.0 58.2 $\pm$ 0.1 62.0 $\pm$ 0.7 EuroSAT-S1 OA 81.7 $\pm$ 0.7 81.5 $\pm$ 0.9 75.4 $\pm$ 0.4 83.6 $\pm$ 0.1 83.9 $\pm$ 0.1 81.7 $\pm$ 0.1 87.2 $\pm$ 0.1 EuroSAT-S2 OA 97.5 $\pm$ 0.0 97.6 $\pm$ 0.1 92.5 $\pm$ 0.1 96.7 $\pm$ 0.0 97.0 $\pm$ 0.1 97.2 $\pm$ 0.1 97.9 $\pm$ 0.1 BigEarthNet-S1 mAP 71.5 $\pm$ 0.4 70.6 $\pm$ 0.4 63.8 $\pm$ 0.1 78.7 $\pm$ 0.0 70.8 $\pm$ 0.0 70.5 $\pm$ 0.0 77.9 $\pm$ 0.0 BigEarthNet-S2 mAP 79.5 $\pm$ 0.5 80.1 $\pm$ 0.1 71.6 $\pm$ 0.1 83.6 $\pm$ 0.0 76.4 $\pm$ 0.0 75.5 $\pm$ 0.0 79.0 $\pm$ 0.0 LC100Cls-S3 mAP 91.3 $\pm$ 0.3 91.4 $\pm$ 0.5 88.9 $\pm$ 0.1 89.5 $\pm$ 0.0 93.3 $\pm$ 0.4 DFC2020-S1 mIoU 49.9 $\pm$ 0.4 50.8 $\pm$ 0.5 45.4 $\pm$ 0.1 52.8 $\pm$ 0.6 52.7 $\pm$ 0.1 49.7 $\pm$ 0.1 52.4 $\pm$ 0.1 DFC2020-S2 mIoU 65.3 $\pm$ 0.6 66.2 $\pm$ 0.7 62.3 $\pm$ 0.0 64.1 $\pm$ 0.3 66.5 $\pm$ 0.0 61.8 $\pm$ 0.1 64.5 $\pm$ 0.1 LC100Seg-S3 mIoU 20.1 $\pm$ 0.4 19.3 $\pm$ 0.5 18.2 $\pm$ 0.1 16.5 $\pm$ 0.1 24.1 $\pm$ 0.0 Flood-S1 mIoU 78.0 $\pm$ 0.1 78.3 $\pm$ 0.3 75.1 $\pm$ 0.1 77.2 $\pm$ 0.1 77.4 $\pm$ 0.1 76.0 $\pm$ 0.1 77.7 $\pm$ 0.0 LCZ-S2 OA 86.6 $\pm$ 0.7 85.3 $\pm$ 0.8 77.4 $\pm$ 0.1 83.6 $\pm$ 0.2 84.1 $\pm$ 0.0 83.0 $\pm$ 0.3 84.4 $\pm$ 0.0 Biomass-S3 RMSE $\downarrow$ 68.1 $\pm$ 0.3 68.3 $\pm$ 0.4 68.7 $\pm$ 0.5 74.1 $\pm$ 0.1 66.3 $\pm$ 0.1 AQ-NO2-S5P RMSE $\downarrow$ 3.4 $\pm$ 0.0 3.4 $\pm$ 0.0 3.4 $\pm$ 0.0 3.3 $\pm$ 0.0 2.8 $\pm$ 0.0 AQ-O3-S5P RMSE $\downarrow$ 1781.3 $\pm$ 29.8 1766.8 $\pm$ 22.1 1741.6 $\pm$ 11.5 1755.6 $\pm$ 19.8 789.4 $\pm$ 2.6